# Deep RL Needs Deep Behavior Analysis: Exploring Implicit Planning by Model-Free Agents in Open-Ended Environments

### v1

@@ 1 Introduction
As reinforcement learning (RL) researchers pursue more complex, naturalistic, and open-ended tasks [1], there is a growing need for tools to understand both the detailed behaviors and the neural dynamics of deep RL agents. This challenge mirrors long-standing problems in neuroscience, where researchers have developed extensive methods for joint behavioral and neural analysis in animals to understand biological intelligence. While ML has adopted neural representation analysis tools from neuroscience, mature behavioral analysis methods from neuroscience, cognitive science, and ethology [2, 3] remain underused in DRL. Open-ended RL tasks typically involve partial observability, memory, planning, and spatial reasoning—yet we still lack a systematic understanding of the strategies agents use to solve them [4, 5, 6, 7, 8]. Further, despite thousands of benchmarks—from video games [9] to 2D/3D gridworlds [10, 11]—there are no standard methods for analyzing or comparing agent behavior across tasks. RL evaluation still focuses primarily on reward curves and aggregate performance, which, while useful, offer limited insight into how agents solve structured tasks. The lack of behavioral analysis is a key bottleneck—limiting insight into why agents succeed or fail. Without understanding behavior, it’s difficult to diagnose algorithmic performance. Applied RL also stands to benefit from removing this bottleneck: more complex tasks require both better agent planning and deeper diagnostics to verify expected behavior [12, 13, 14].
@@
Notably, this is not just a machine learning problem—interpreting behavior in partially observable, reward-driven environments is also a core challenge in neuroscience and cognitive science. Inspired by how animals are studied in neuroscience and ethology, we propose a behavior-first approach to analyzing DRL agents. This approach enables richer insight into both learned behavior and internal representations, drawing on joint behavioral-neural analysis tools that have proven essential for understanding biological intelligence. We ground this proposal in ForageWorld, a novel task suite in which agents forage and survive in procedurally generated, large-scale, partially observable environments with temporally extended dynamics. These pressures reflect the ecological tradeoffs animals encounter during natural foraging [15, 16]. ForageWorld builds on Craftax [11], combining complex naturalistic environment features with efficient training. We extend Craftax with features such as patchy resources and structured threats to better capture key elements of animal foraging. These additions place strong demands on memory, spatial reasoning, and long-term planning, presenting a non-trivial challenge for modern DRL agents. In addition, foraging serves as a useful metaphor for core robotics tasks such as navigation, cleaning, and search and rescue [17], and underlies many popular RL benchmarks, including Minecraft and NetHack [18].
@@
To support behavior-first evaluation in DRL, we introduce an analysis framework (Table 1) that maps key behavioral and neural targets to diagnostic methods. This framework captures the core dimensions analyzed in this work and is reusable across neuroAI, interpretability, and broader RL applications.
@@ 2.1 Open-Ended RL and the Need for Behavioral Analysis
Generalization in RL refers to the ability to transfer knowledge to new environments or tasks, and is critical for both practical applications and theoretical understanding in AI [19]. However, many RL algorithms struggle to meaningfully generalize beyond their training tasks [20, 21]. This bottleneck has renewed interest in open-ended RL tasks that require rich exploration, subgoal discovery, and long-horizon reasoning [1, 22]. Open-ended RL tasks typically (i) scale with the learner, requiring skill composition in large state spaces, and (ii) incorporate naturalistic structure to better reflect real-world environments [22]. Many recent benchmarks embody this open-ended perspective—emphasizing compositional skills, dynamic goals, procedural generation, and expansive state spaces[1, 10, 11, 12, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32].
@@ 2.2 Joint Behavioral and Neural Analyses for Artificial Agents
To date, open-ended RL benchmarks have not drawn sustained interest from neuroscience or cognitive science, leading to a lack of deeper analysis frameworks in this research area. One reason is that few open-ended tasks reflect how animals explore, plan, and decide [3, 33]. Thus, there is a need for more naturalistic benchmarks—and for analysis pipelines that study agents with the same rigor applied to animal behavior and neural computation. Recent neuroscience and cognitive science work has primarily applied brain-inspired tools to ML systems for comparing internal representations [34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. This lack of behavioral emphasis mirrors limitations in how internal dynamics are typically studied—often in isolation from behavior, or without the tools needed to reveal structure. To date, most of this work has focused on representational similarity, not on joint analysis of behavior and underlying computation. A few recent exceptions—mostly outside RL—have begun to explore this connection [40, 44, 45, 12, 46].
Despite this challenge, a key lesson from neuroscience is that neural activity alone is often insufficient to explain computation [3, 47, 48, 49, 50]. For example, recent findings show that common similarity metrics can yield inconsistent conclusions [41, 51]. Several recent works argue that neuroAI should prioritize aligning behavior and internal dynamics in naturalistic settings, rather than relying on coarse similarity metrics in toy tasks [52, 50]. We thus adopt a behavior-first approach to studying DRL agents in the naturalistic, open-ended task ForageWorld, characterizing their strategies and internal representations using neuroscience-inspired tools—as is done in biological systems.
Figure 1: Model-free agents exhibit structured exploration and revisitation behavior in an open-ended, partially observable environment. (A) A 9$\times$11 grid view (left, local observation window) shows the agent’s local observation, positioned within the full 96$\times$96 environment (right). Cows (food) diffuse from spawn points and deplete when consumed; lakes (drink) are fixed and unlimited; predators intermittently pursue agents in view. (B) Agent observations include the visual window, inventory, and internal states (e.g., health, hunger, fatigue). This agent is sleeping—typically in corners to reduce predator contact. (C) Trajectories from a single episode: three early sequential exploratory paths (left), initial revisitation (middle), and full path with revisited patches in orange (right).
@@ 3 Methods
This section summarizes the environment, model architectures, training procedures, and analysis framework. Full code is available at https://anonymous.4open.science/r/ForageWorld/, with additional details in the appendix.
@@ 3.1 ForageWorld: Naturalistic Foraging with Internal States
We build on the Craftax benchmark [11], modifying it to support more naturalistic behavior under partial observability. ForageWorld introduces ecological constraints and uncertainty to better reflect real-world cognitive demands and enable deeper analysis. In particular, the need to localize food and avoid predators increases the memory demands relative to the original Craftax task. Cows have an arena-wide limit, spawn at fixed locations, and diffuse over time—leading to temporary, localized food depletion and requiring the agent to leave and revisit patches periodically (Figure 1).
Drinking from procedurally placed water sources maintains hydration. Predators intermittently appear near food patches and pursue agents when within line of sight (Figure 1). Agents manage hunger, thirst, fatigue, and health. Fatigue recovers only through sleep, starvation and extreme thirst reduce health, and death occurs if the health variable hits zero. Sleep can only be performed when energy is below 50% of maximum. It immobilizes the agent while gradually restoring energy and, if applicable, health (Figure 1). Arenas are procedurally generated with randomized layouts, ensuring per-episode novelty. Unlike Craftax, our reward function is focused on agent survival: to receive reward, the agent must maintain its physiological resources (health, food, drink, energy) above half their maximum values. This objective correlates strongly with survival time and encourages the agent to maintain a buffer in resource levels—without promoting overconsumption or maxing out all variables. Specifically, the reward function is:
@@
As defined in Equation 1, the agent’s reward is a function of physiological variables: it receives positive reward only when all key survival resources—health, food, drink, and energy—are maintained above a threshold.
@@ 3.2 RL Architectures and Model Variants
Our primary architecture was PPO with a GRU-based recurrent core (PPO-RNN), detailed in subsection A.6. In some experiments, we added an auxiliary objective to encourage interpretable spatial representations. In these experiments, the model also outputs its predicted position $(x_{t},y_{t})$ relative to the first-timestep origin, via a small fully connected head that shares the RNN hidden state $h_{t}$ owith the policy and value heads. We train this auxiliary objective jointly with the PPO loss, using an L2 regression loss:
@@
To introduce biologically-inspired sparsity, we prune the weights of the RNN core using JaxPruner during training [53]. We targeted 90% sparsity across model layers, approximating biological brain sparsity, which ranges from 70% to 94% across species [54, 55]. We used magnitude-based pruning. Pruned agents performed comparably to unpruned ones and, in some cases, showed improved decoding of spatial variables from hidden states (Figure 16)–consistent with findings that sparse connectivity improves modularity and interpretability [56, 57, 58].
@@ 3.3 Decoding Allocentric Position and Planning Horizons
To evaluate memory and planning, we test whether hidden states encode past and future spatial positions. If location is internally represented, it may enable both revisitation and planning. To measure this, we first record hidden states $h_{t}$ and ground-truth positions $(x_{t},y_{t})$ throughout each episode. Then, at each timestep $t$, we train a decoder to predict allocentric displacement $Y_{t+\Delta t}=(\Delta x,\Delta y)$ between the current hidden state $h_{t}$ and the agent’s position $\Delta t$ steps in the past or future, using only $h_{t}\in\mathbb{R}^{512}$. This method tests whether the RNN encodes allocentric spatial information at various planning and memory horizons—analogous to representations in animal navigation systems. Additionally, we train separate decoders for allocentric (relative to origin) and egocentric (relative to body orientation) coordinates to assess which coordinate frame the agent uses internally. We use ridge regression as the decoding model to preserve interpretability. Decoders were trained on the first 75% of each episode and evaluated on the final 25% to assess generalization. Accuracy is reported as RMSE, benchmarked against an average displacement-based baseline. All behavioral analyses were conducted in held-out test arenas, with weights frozen throughout, to ensure that behavior reflects generalization rather than ongoing learning. See subsection C.2 and subsection C.3 for full training details.
@@ 4.1 Structured Exploration and Revisitation in Novel Environments
We begin by examining how agents explore novel environments and transition to revisitation strategies. Foraging has long served as a rich behavioral paradigm for studying goal-directed exploration in animal neuroscience [59, 60]. In rodents, foraging in novel, partially observable environments unfolds in distinct behavioral phases: mice initially engage in broad exploration, followed by direct and efficient revisitation of previously discovered locations once the environment is mapped [61].
Using trajectory visualizations, behavioral segmentation, and spatial entropy metrics, we find that ForageWorld induces structured, phase-like foraging behaviors in trained DRL agents—analogous to those observed in biological systems. Standard performance metrics confirm that agents are competent at the task, showing long survival times and balanced resource levels over training (Figure 4). However, our behavioral analysis reveals that DRL agents equipped only with RNNs—i.e., without explicit world models—develop in-context adaptive learning and exploration strategies that evolve across episodes.
Strikingly, after initially mapping a novel arena, agents shift into revisitation behaviors that reflect structured, multi-objective decision-making—mirroring phase transitions in animal foraging [61] and information seeking [62, 63, 64]. Their early exploration trajectories qualitatively resemble known patterns of insect search behavior (Figure 8). Agents generated outwardly spiraling, azimuthally rotating loops from their starting position that gradually expanded to cover most of the arena (Figure 1). Even during revisitation phases, exploration persisted—most successful agents had covered the majority of the arena by the end of each episode (Figure 6). These extended trajectories also supported predator avoidance, enabling agents to discover shortcut paths between safe regions and resource patches (Figure 7). Similar patterns have been observed in ants and bees during nest displacement or early flights from the hive (Figure 8).
@@ 4.2 Multi-Objective Foraging and Strategic Patch Use
Having characterized agents’ exploration and revisitation behaviors, we next examine how they manage and select among remembered locations. After initial exploration, agents displayed a clear shift to revisitation—returning to remembered patches via direct routes. Similar to rodents navigating partially observable mazes [61], our agents exhibit a rapid transition from broad exploration to targeted revisitation (Figure 1). This transition reflects phase-structured foraging as seen in rodents, while the spatial patterns resemble insect search behavior (Figure 8). Agents also periodically return to their starting position—potentially a reorientation strategy in the absence of an explicit spatial map.
@@
Figure 2: Performance metrics and model ablations motivate deeper behavioral-neural analysis. (Top left) ForageWorld requires memory and substantial network capacity to learn. Replacing the 512-unit RNN with a feedforward network significantly impairs performance, as does downsizing to a 64-unit network ( 550k parameters). A midsize 128-unit network also underperforms when pruned, suggesting that high capacity is needed to support sparsity. (Top right) Pruning does not degrade training performance but improves the spatial interpretability of internal representations (see Figure 16). (Bottom left) Removing the auxiliary path integration loss reduces performance in large arenas—but not small ones. This is not evident from reward curves or predator milestones (see Figure 11), but is explained by the model’s inability to predict its position, which appears critical for large-scale navigation (see Figure 17). (Bottom right) Limiting perception to a forward-facing field of view improves early learning—perhaps due to reduced input complexity—but final performance conceals major behavioral differences (see Figure 15). (All plots) Curves show the time-weighted EMA mean across 5 random seeds; shaded regions show one standard deviation.
@@
Agents are not given an explicit list of patch locations. Instead, revisitation behavior emerges from recurrent state dynamics encoded during exploration and is reconstructed for analysis from logged trajectories. This result shows that model-free RNN agents can develop memory-guided, goal-directed behavior—such as selective revisitation—without explicit world models or symbolic memory structures. This memory dependence is further supported by a substantial performance gap between PPO agents with and without recurrence (Figure 2, Figure 10). Like animals that balance intake against physiological constraints (e.g., stomach volume, satiation), our agents are rewarded for maintaining low hunger, thirst, and fatigue—rather than for rapidly depleting resource patches. Under our baseline task parameters, expert agents ate at an average rate of 0.011 (approximately once every 95 steps), drank at 0.034, and slept at 0.262. This emergent spacing enabled us to test whether agents made memory-guided, multi-objective decisions when selecting which patches to revisit. We analyzed patch revisitation choices and found that agents integrated spatial and task-relevant features in a multi-objective manner (Figure 3, Figure 7).
@@
Figure 3: Memory-guided, multi-objective revisitation strategies emerge in model-free agents without world models. Generalized linear model (GLM) coefficients for patch history variables predicting an agent’s choice to revisit one patch over others. Choices are defined 50 timesteps before each patch eat event. From left to right: agents prefer patches with fewer prior eat actions (EatRate); show no preference for water proximity (DrinkRate); avoid patches with more predator encounters (PredRate); prefer patches visited more recently (Recency); show mild preference for longer dwell time (Dwelltime); prefer patches with more observed cows (CowCount); and prefer patches with higher prior position prediction error (Uncertainty). Significance levels: * $<$ 0.05, ** $<$ 0.01, *** $<$ 0.001. See Figure 13 for full model output and VIF analysis.

### v2

@@ 1 Introduction
As reinforcement learning (RL) researchers pursue more complex, naturalistic, and open-ended tasks [team2021open], there is a growing need for tools to understand both the detailed behaviors and the neural dynamics of deep RL agents [bastankhah2025demystifying, liu2024single]. This challenge mirrors long-standing problems in neuroscience, where researchers have developed extensive methods for joint behavioral and neural analysis in animals to understand biological intelligence [pedamonti2025hippocampus]. While machine learning (ML) has adopted neural representation analysis tools from neuroscience, mature behavioral analysis methods from neuroscience, cognitive science, and ethology [krakauer2017neuroscience, niv2021primacy] remain underused in DRL. Open-ended RL tasks typically involve partial observability, memory, planning, and spatial reasoning—yet we still lack a systematic understanding of the strategies agents use to solve them [wolbers2014challenges, epstein2017cognitive, ekstrom2017interacting, zhu2021deep, parra2023map, tamir2025]. Further, despite thousands of benchmarks—from video games [mnih2015human] to 2D/3D gridworlds [chevalier2023minigrid, matthews2024craftax]—there are no standard methods for analyzing or comparing agent behavior across tasks. RL evaluation still focuses primarily on reward curves and aggregate performance, which, while useful, offer limited insight into how agents solve structured tasks. The lack of behavioral analysis is a key bottleneck—limiting insight into why agents succeed or fail. Without understanding behavior, it’s difficult to diagnose algorithmic performance. Applied RL also stands to benefit from removing this bottleneck: more complex tasks require both better agent planning and deeper diagnostics to verify expected behavior [chung2023thinker, fetsch2016importance, fu2021minimizing].
@@
Notably, this is not just a machine learning problem—interpreting behavior in partially observable, reward-driven environments is also a core challenge in neuroscience and cognitive science. Inspired by how animals are studied in neuroscience and ethology, we propose a behavior-first approach to analyzing DRL agents. This approach enables richer insight into both learned behavior and internal representations, drawing on joint behavioral-neural analysis tools that have proven essential for understanding biological intelligence. We ground this proposal in ForageWorld, a novel task suite in which agents forage and survive in procedurally generated, large-scale, partially observable environments with temporally extended dynamics. These pressures reflect the ecological tradeoffs animals encounter during natural foraging [kolling2014multiple, lai2024robot], and align with calls to push foraging tasks in neuroAI beyond the traditional bandit paradigms [hall2019revisiting, morimoto2019foraging, boccara2019entorhinal, badman2020multiscale, jiang2022hippocampal, grima2025foraging]. ForageWorld builds on Craftax [matthews2024craftax], combining complex naturalistic environment features with efficient training. We extend the GPU-accelerated Craftax with features such as patchy resources and structured threats to better capture key elements of animal foraging. These additions place strong demands on memory, spatial reasoning, and long-term planning, presenting a non-trivial challenge for modern DRL agents. In addition, foraging serves as a useful metaphor for core robotics tasks such as navigation, cleaning, and search and rescue [Winfield2009], and underlies many popular RL benchmarks, including Minecraft and NetHack [fan2022minedojo].
@@
To support behavior-first evaluation in DRL, we introduce an analysis framework (Table 1) that maps key behavioral and neural targets to diagnostic methods. This framework is showcased by the results in this work and is reusable across neuroAI, interpretability, and broader RL applications.
@@ 2.1 Open-Ended RL and the Need for Behavioral Analysis
Generalization in RL refers to the ability to transfer knowledge to new environments or tasks, and is critical for both practical applications and theoretical understanding in AI [sutton1999reinforcement]. However, many RL algorithms struggle to generalize beyond their training tasks [packer2018assessing, cobbe2019quantifying]. This bottleneck has renewed interest in open-ended RL tasks that require rich exploration, subgoal discovery, and long-horizon reasoning [team2021open, jiang2023general]. Open-ended RL tasks typically (i) scale with the learner, requiring skill composition in large state spaces, and (ii) incorporate naturalistic structure to better reflect real-world environments [jiang2023general]. Many recent benchmarks embody this open-ended perspective—emphasizing compositional skills, dynamic goals, procedural generation, and expansive state spaces[team2021open, chevalier2023minigrid, matthews2024craftax, chung2023thinker, wei2022honor, norstein2022open, cai2023open, samvelyan2023maestro, zhang2023creative, tomilin2023coom, yuan2023skill, jucys2024, nikulin2024xland, davidson2025goals].
@@ 2.2 Joint Behavioral and Neural Analyses for Artificial Agents
To date, open-ended RL benchmarks have not drawn sustained interest from neuroscience or cognitive science, leading to a lack of deeper analysis frameworks in this research area. One reason is that few open-ended tasks reflect how animals explore, plan, and decide [niv2021primacy, wise2024naturalistic]. Thus, there is a need for more naturalistic benchmarks—and for analysis pipelines that study agents with the same rigor applied to animal behavior and neural computation. Recent neuroscience and cognitive science work has primarily applied brain-inspired tools to ML systems for comparing internal representations [schrimpf2018brain, richards2019deep, schrimpf2020integrative, nonaka2021brain, saxe2021if, ostrow2023beyond, momennejad2023evaluating, soni2024conclusions, wang2024investigating, carvalho2024predictive]. This lack of behavioral emphasis mirrors limitations in how internal dynamics are typically studied—often in isolation from behavior, or without the tools needed to reveal structure. To date, most of this work has focused on representational similarity, not on joint analysis of behavior and underlying neural computation. A few recent exceptions—mostly outside RL—have begun to explore interpretability with a greater focus on behavior [momennejad2023evaluating, yesiltepe2023entropy, hwang2023grid, chung2023thinker, schneider2023learnable]. Researchers have also made calls for more interpretability work in the RL domain as well [abney2025advancing, verma2018programmatically, cheng2025survey].
Despite this challenge, a key lesson from neuroscience is that neural activity alone is often insufficient to explain computation [niv2021primacy, hirsch1986nothing, jonas2017could, geirhos2023don, linsley2025AIbrain]. For example, recent findings show that common similarity metrics can yield inconsistent conclusions [soni2024conclusions, cui2022deconfounded]. Several recent works argue that neuroAI should prioritize aligning behavior and internal dynamics in naturalistic settings, rather than relying on coarse similarity metrics in toy tasks [cao2024explanatory, linsley2025AIbrain]. We thus adopt a behavior-first approach to studying DRL agents in the naturalistic, open-ended task ForageWorld, characterizing their strategies and internal representations using neuroscience-inspired tools—as is done in biological systems.
[…]
@@ 3 Methods
This section summarizes the environment, model architectures, training procedures, and analysis framework we used in this work. Full code for ForageWorld and our analysis is available at https://github.com/RileySE/Craftax-Foraging/tree/foraging, with additional details in the appendix.
@@ 3.1 ForageWorld: Naturalistic Foraging with Internal States
We build on the Craftax benchmark [matthews2024craftax], modifying it to support more naturalistic behavior under partial observability. ForageWorld introduces ecological constraints and uncertainty to better reflect real-world cognitive demands and enable deeper analysis. In particular, the need to localize food and avoid predators increases the memory demands relative to the original Craftax task. Cows have an arena-wide count limit, spawn at fixed locations, and diffuse over time—leading to temporary, localized food depletion and requiring the agent to leave and revisit patches periodically (Figure 1).
Drinking from procedurally placed water sources maintains hydration. Predators intermittently appear near food patches and pursue agents when within line of sight (Figure 1). Agents manage hunger, thirst, fatigue, and health. Fatigue recovers only through sleep, starvation and extreme thirst reduce health, and death occurs if the health variable hits zero. Sleep can only be performed when energy is below 50% of maximum. It immobilizes the agent while gradually restoring energy and, if applicable, health (Figure 1). Arenas are procedurally generated with randomized layouts, ensuring per-episode novelty. Unlike Craftax, our reward function is focused on agent survival: to receive reward, the agent must maintain its physiological resources (health, food, drink, energy) above half their maximum values. This objective correlates strongly with survival time and encourages the agent to maintain a buffer in resource levels—without promoting overconsumption or maxing out all variables [leeevaaa]. Specifically, the reward function is:
@@
As defined in Equation 1, the agent’s reward is a function of physiological variables: it receives increasing positive reward for each key survival resource maintained above a threshold—health, food, drink, and energy.
@@ 3.2 RL Architectures and Model Variants
Our primary architecture was PPO with a GRU-based recurrent core (PPO-RNN), detailed in subsection A.6. In some experiments, we added an auxiliary objective to encourage interpretable spatial representations. In these experiments, the model also outputs its predicted position $(x_{t},y_{t})$ relative to the first-timestep origin (path integration), via a small fully connected head that shares the RNN hidden state $h_{t}$ with the policy and value heads. We train this auxiliary objective jointly with the PPO loss, using an L2 regression loss:
@@
To introduce biologically-inspired sparsity, we prune the weights of the RNN core using JaxPruner during training [lee2024jaxpruner]. We targeted 90% sparsity across model layers, approximating biological brain sparsity, which ranges from 70% to 94% across species [broido2019scale, zhang2023universal]. We used magnitude-based pruning. Pruned agents performed comparably to unpruned ones and, in some cases, showed improved decoding of spatial variables from hidden states (Figure 16)–consistent with findings that sparse connectivity improves modularity and interpretability [frankle2019lottery, michel2019sixteen, tumma2023leveraging].
@@ 3.3 Decoding Allocentric Position and Planning Horizons
To evaluate memory and planning, we test whether hidden states encode past and future spatial positions. If location is internally represented, it may enable revisitation and planning. We first record hidden states $h_{t}$ and ground-truth positions $(x_{t},y_{t})$ throughout each episode. Then, at each timestep $t$, we train a decoder to predict allocentric displacement $Y_{t+\Delta t}=(\Delta x,\Delta y)$ between the current hidden state $h_{t}$ and the agent’s position $\Delta t$ steps in the past or future, using only $h_{t}\in\mathbb{R}^{512}$. This method tests whether the RNN encodes allocentric spatial information at various planning and memory horizons—analogous to representations in animal navigation systems. Additionally, we train separate decoders for allocentric (relative to origin) and egocentric (relative to body orientation) coordinates to assess which coordinate frame the agent uses internally. We use ridge regression as the decoding model to preserve interpretability. Decoders were trained on the first 75% of each episode and evaluated on the final 25% to assess generalization. Accuracy is reported as RMSE, benchmarked against an average displacement-based baseline. All behavioral analyses were conducted in held-out test arenas, with weights frozen throughout, to ensure that behavior reflects generalization rather than ongoing learning. See subsection C.2 and subsection C.3 for full training details.
@@ 4.1 Structured Exploration and Revisitation in Novel Environments
We begin by examining how agents explore novel environments and transition to revisitation strategies. Foraging has long served as a rich behavioral paradigm for studying goal-directed exploration in animal neuroscience [hayden2018economic, hills2006animal]. In rodents, foraging in novel, partially observable environments unfolds in distinct behavioral phases: mice initially engage in broad exploration, followed by direct and efficient revisitation of previously discovered locations once the environment is mapped [rosenberg2021mice].
Using trajectory visualizations, behavioral segmentation, and spatial entropy metrics, we find that ForageWorld induces structured, phase-like foraging behaviors in trained DRL agents—analogous to those observed in biological systems. Standard performance metrics confirm that agents are competent at the task, showing long survival times (directly proportional to return) after training (Figure 2). However, our behavioral analysis reveals that DRL agents equipped only with RNNs—i.e., without explicit world models—develop in-context adaptive learning and exploration strategies that evolve across episodes.
Strikingly, after initially mapping a novel arena, agents shift into revisitation behaviors that reflect structured, multi-objective decision-making—mirroring phase transitions in animal foraging [rosenberg2021mice] and information seeking [oudeyer2016intrinsic, bussell2023representations, monosov2024curiosity]. Their early exploration trajectories qualitatively resemble known patterns of insect search behavior (Figure 8). Agents generated outwardly spiraling, azimuthally rotating loops from their starting position that gradually expanded to cover most of the arena (Figure 1). Even during revisitation phases, exploration persisted—most successful agents had covered the majority of the arena by the end of each episode (Figure 6). These extended trajectories also supported predator avoidance, enabling agents to discover shortcut paths between safe regions and resource patches (Figure 7). Similar patterns have been observed in ants and bees during nest displacement or early flights from the hive (Figure 8).
@@ 4.2 Multi-Objective Foraging and Strategic Patch Use
Having characterized agents’ exploration and revisitation behaviors, we next examine how they select among remembered locations. After initial exploration, agents displayed a clear shift to revisitation—returning to remembered patches via direct routes. Similar to rodents navigating partially observable mazes [rosenberg2021mice], our agents exhibit a rapid transition from broad exploration to targeted revisitation (Figure 1). This transition mimics phase-structured foraging as seen in rodents, while exploration resembles insect search behavior (Figure 8). Agents also periodically return to their starting position—potentially a reorientation strategy in the absence of an explicit spatial map.
@@
Figure 2: Performance metrics and model ablations motivate deeper behavioral-neural analysis. (Top left) ForageWorld requires memory and substantial network capacity to learn. Replacing the 512-unit RNN with a feedforward network significantly impairs performance, as does downsizing to a 64-unit network (550k parameters). A midsize 128-unit network also underperforms when pruned, suggesting that high capacity is needed to support sparsity. (Top right) Pruning does not degrade training performance but improves the spatial interpretability of internal representations (see Figure 16). (Bottom left) Removing the auxiliary path integration loss reduces performance in large arenas—but not small ones. Thus, the ability to predict current self-position appears critical for larger-scale navigation (see Figure 17 for representational effects). (Bottom right) Limiting perception to a forward-facing field of view improves early learning—perhaps due to reduced input complexity—but final performance conceals behavioral differences (see Figure 15). (All plots) Curves show the time-weighted EMA mean across 5 random seeds; shaded regions show one standard deviation.
@@
Agents are not given an explicit list of patch locations. Instead, revisitation behavior emerges from recurrent state dynamics encoded during exploration and is reconstructed for analysis from logged trajectories. This result shows that model-free RNN agents can develop memory-guided, goal-directed behavior—such as selective revisitation—without explicit world models or symbolic memory structures. This memory dependence is further supported by a substantial performance gap between PPO agents with and without recurrence (Figure 2, Figure 11, Figure 10). Like animals that balance intake against physiological constraints (e.g., stomach volume, satiation), our agents are rewarded for maintaining low hunger, thirst, and fatigue—rather than for rapidly depleting resource patches. Under our baseline task parameters, expert agents ate at an average rate of 0.011 (approximately once every 95 steps), drank at 0.034, and slept at 0.262. This emergent spacing enabled us to test whether agents made memory-guided, multi-objective decisions when selecting which patches to revisit. We analyzed patch revisitation choices and found that agents integrated spatial and task-relevant features in a multi-objective manner (Figure 3, Figure 7).
@@
Figure 3: Memory-guided, multi-objective revisitation strategies emerge in model-free agents without world models. Generalized linear model (GLM) coefficients for patch history variables predicting an agent’s choice to revisit one patch over others. Choices are defined 50 timesteps before each patch eat event. From left to right: agents prefer patches with fewer prior eat actions (EatRate); show no preference for water proximity (DrinkRate); avoid patches with more predator encounters (PredRate); prefer patches visited more recently (Recency); show mild preference for longer dwell time (Dwelltime); prefer patches with more observed cows (CowCount); and prefer patches with higher prior position prediction error (Uncertainty). Significance levels: * << 0.05, ** << 0.01, *** << 0.001. Figure 13 has model output and VIF analysis. Error bars are 95% confidence intervals (CI).